Grade 10Mechanics

Time period of an earth satellite is 83 minutes. Time period of another satellite at a distance of three times earths radius is

Sanjana
8 Years agoGrade 10

1 Answer

Khimraj
8 Years ago
for first satellite r =R
and fot second r = 4R
then
T2/T1 = (r2/r1)3/2 
T2 = T1*43/2 = 83*43/2 = 664 minutes